# Part I: Nonlinear Codes

### v1

@@ Abstract
A de Bruijn arrays code is a set of $r\times s$ binary doubly-periodic arrays such that each binary $n\times m$ matrix is contained exactly once as a window in one of the arrays. Such a set of arrays can be viewed as a two-dimensional generalization of a perfect factor in the de Bruijn graph. Necessary conditions for the existence of such arrays are given. Several direct constructions and recursive constructions for such arrays are given. A framework for a theory of two-dimensional feedback shift register which is akin to (one-dimensional) feedback shift registers is suggested.
@@ I Introduction
Generalizations of one-dimensional sequences and codes to higher dimensions is quite natural from both theoretical and practical point of view. Such generalizations were considered for various structures such as error-correcting codes [1, 47], burst-correcting codes [2, 3, 15, 16], constrained codes [50, 54], and de Bruijn sequences [8, 34, 42]. This paper considers a generalization of one-dimensional sequences with a window property to two-dimensional arrays with a window property. For simplicity in this paper only binary arrays and usually binary sequences are considered, although some of the results can be generalized to any alphabet size and some sequences over larger alphabets can be used to construct binary arrays.
@@
A de Bruijn array (or a perfect map) is an $r\times s$ doubly-periodic array, such that each $n\times m$ binary matrix appears exactly once as a window in the array. Such an array will be referred to as $(r,s;n,m)$-DBA.
Perfect maps were first presented by Reed and Stewart [46] and later in Gordon [21] and Clapham [6]. A very simple construction based on a span $n$ de Bruijn sequence and a de Bruijn sequence over an alphabet of size $2^{n}$ was presented by Ma [31]. This construction will be modified later in one of our constructions. Fan, Fan, Ma, and Siu [17] gave a two-dimensional graphical representation for perfect maps. They also presented two recursive constructions for such de Bruijn arrays. An important construction for perfect maps and another graphical representation was given in [8]. Finally, Paterson [42] proved that the necessary conditions for the existence of de Bruijn arrays are also sufficient. These arrays have found a variety of applications. They were used in pattern recognition for structured light systems as described in Geng [19], Lin, Nie, and Song [29], Morano, Ozturk, Conn, Dubin, Zietz, and Nissanov [40], Salvi, Fernandez, Pribanic, and Llado [48], and Salvi, Pagès, and Batlle [49]. They are also used in transferring planar surface into a sensitive touch screen display, see Dai and Chung [7], in camera localization as described by Szentandrasi, Zachariás̆, Havel, Herout, Dubska, and Kajan [53], and in one-shot shape acquisition, e.g., Pagès, Salvi, Collewet, and Forest [41]. Finally, they can be applied in surface measurements as described in Kiyasu, Hoshino, Yano, and Fujimura [26] and in Spoelder, Vos, Petriu, and Groen [52], and also in coded aperture imaging as was presented for example in Gottesman and Fenimore [22].
The are of de Bruijn arrays (which were generally called perfect maps) were extensively covered in the literature. Some more work on this topic and related structures was done by [23, 24, 25, 30, 32, 33, 34, 35, 36, 37]
The goal of the current paper is to extend the results on de Bruijn arrays into de Bruijn arrays codes defined as follows.
@@
A de Bruijn arrays code is a set of $r\times s$ doubly-periodic arrays, such that each nonzero $n\times m$ matrix appears exactly once as a window in one of the arrays. Such a set of arrays will be referred to as $(r,s;n,m)$-DBAC.
@@
Lemma 1 implies that $r$, $s$, and $\Delta$, are powers of two. Henceforth, we use $2^{k}$ instead of $r$ and $2^{t}$ instead of $s$. One of the goals in the research on DBACs is to prove that the necessary conditions of Lemma 1 are sufficient. In the discussion which is given in this paper it will be also shown how the definition of de Bruijn arrays codes generalizes some one-dimensional definitions.
@@
The rest of this paper is organized as follows. In Section II we present the basic definitions and theory for the arrays which will be considered in this work. The section consists of five subsections. To understand the theory we have to understand some basic concepts associated with shift registers and their sequences. These will be described in Section II-A. Some state diagrams of these registers are the one-dimensional analogs of the de Bruijn arrays codes. Two well-known operators defined for the one-dimensional theory of shift registers will play an important role in our exposition, the shift operator ${\bf E}$, and the ${\bf D}$-morphism. The shift operator ${\bf E}$ will be discussed in Section II-B and the emphasis will be on its role in defining sequences which are generated by linear feedback shift registers. It will be also important in the definition of linear complexity of sequences which will be used for some of the constructions of the two-dimensional arrays. The operator ${\bf D}$ will be discussed in Section II-C. This operator defines a homomorphism between de Bruijn graphs of different orders. It will be used in the recursive constructions of de Bruijn arrays and it is strongly connected to the shift operator ${\bf E}$. The structures that will be heavily used in both the direct constructions and the recursive constructions are the perfect factors which will be discussed in Section II-D. Perfect factors were used in the constructions of de Bruijn arrays and will be used in the constructions of de Bruijn arrays codes. They will be used in the direct constructions and in the recursive constructions. In the recursive constructions they will be used as columns of the arrays and also as sequences which define how the constructions are applied with the inverse of the operator ${\bf D}$. The one-dimensional sequences are associated with the de Bruijn graph and for the two-dimensional arrays two different graph representations will be presented in Section II-E. In particular we will examine how one-dimensional nonsingular feedback shift registers are generalized into two-dimensions. Many constructions will be presented in the paper, some of these constructions are direct constructions and some of them are recursive constructions. Section III is devoted to five direct constructions. One construction is obtained by ordering the sequences of a perfect factor. In a second construction the sequences of a perfect factor are ordered in a way that all the possible shifts between two sequences are obtained. A third construction generalizes the construction of Ma [31] using a perfect factor over a large alphabet and a binary de Bruijn sequence. The last two direct constructions are two-dimensional generalizations of two of the most celebrated nonsingular shift registers. These generalizations should be done carefully to obtain a state diagram with cycles of the same length. Five recursive constructions will be presented in Section IV. All these constructions are generalizations of constructions of de Bruijn arrays presented by Fan, Fan, Ma, and Siu [17]. Perfect factors have to be used carefully for these constructions and the chosen arrays on which the constructions are applied should be also chosen in an appropriate way. In Section V another technique to generate $(2^{k},2^{t};n,m)$-DBACs by joining codewords (cycles) of an $(2^{k},2^{t^{\prime}};n,m)$-DBAC with shorter codewords, i.e., $t^{\prime}<t$, is presented. This technique can be applied if all the codewords are cycles of the same length in the same state diagram of the two-dimensional generalization de Bruijn graph. Section VI includes conclusion and future work.
MacWilliams and Sloane [32] gave a construction for another type of arrays defined as follows.
Definition 3
A pseudo-random array ${\cal A}$ is an $r\times s$ doubly-periodic array such that each $n\times m$ nonzero matrix is contained exactly once as a window in the array. Moreover, if ${\cal A}$ is such an array, then ${\cal A}+{\cal A}^{\prime}$, where ${\cal A}^{\prime}$ is a nontrivial shift of ${\cal A}$ and the addition is performed bit-by-bit, is another nontrivial shift of ${\cal A}$. This is a shift-and-add-property. Such an array will be called an $(r,s;n,m)$-PRA.
Definition 4
A shortened perfect map (or a shortened de Bruijn array) is an $r\times s$ binary doubly-periodic array, where ${rs=2^{nm}-1}$, such that each nonzero $n\times m$ matrix appears exactly once as a window in the array. Such an array will be called an $(r,s;n,m)$-SDBA.
The definition of pseudo-random arrays is generalized to pseudo-random arrays codes and it will be considered in the follow-up paper [5]. They were also considered in a preliminary version that discussed these array codes [11]. In the current paper the constructions of DBACs will be based on combinatorial techniques and the constructed codes will be nonlinear. In the follow-up paper [5] the constructions of pseudo-random arrays codes will be based on algebraic techniques and the constructed codes will be linear.
@@ II Preliminaries
This section is devoted to the basic concepts which will be used in our constructions of DBACs and also in explaining how the one-dimensional setup is generalized to the two-dimenstional setup.
@@ II-A Shift-register sequences
A feedback shift register of order $n$ (an FSRn in short) has $2^{n}$ states, represented by the set of the $2^{n}$ binary words of length $n$. The register has $n$ cells, which are binary storage elements, where each cell stores at each stage one of the bits of the current state $x=(x_{1},x_{2},\ldots,x_{n})$. An FSRn is depicted in Fig. 1.
@@
Shift registers and their sequences were extensively studied and two comprehensive books on them are [10, 20]. The known material in this section is presented in these books. Given the word $(x_{1},x_{2},\ldots,x_{n})$ which is a state in the FSRn, $x_{i}$ is stored in the $i$-th cell of the FSRn. The $n$ cells are connected to another logic element which computes a Boolean feedback function $f(x_{1},x_{2},\ldots,x_{n})$. At periodic intervals which are controlled by a master clock, $x_{2}$ is transferred to $x_{1}$, $x_{3}$ to $x_{2}$, and so on until $x_{n}$ is transferred to $x_{n-1}$. The value of the feedback function is transferred to $x_{n}$ and hence it is common to denote $x_{n+1}=f(x_{1},x_{2},\ldots,x_{n})$. The register starts to work with an initial state $(a_{1},a_{2},\ldots,a_{n})$, where $a_{i}$, $1\leq i\leq n$, is the initial value stored in the $i$-th cell. The feedback function $f$ is a Boolean function and hence it can be also represented by a truth table. There are $2^{n}$ possible distinct values for $x_{1},x_{2},\ldots,x_{n}$, i.e., there are $2^{n}$ distinct states and each one can have a value of either 0 or 1. Hence, there are $2^{2^{n}}$ different FSRn, but not all of these functions are of interest.
Each FSRn has a state diagram which is a graph with $2^{n}$ vertices (the states of the FSRn). Given an FSRn with a feedback function $f(x_{1},x_{2},\ldots,x_{n})$, the vertex $(x_{1},x_{2},\ldots,x_{n})$ in the associated state diagram has an edge to the vertex $(x_{2},\ldots,x_{n},x_{n+1})$ if $x_{n+1}=f(x_{1},x_{2},\ldots,x_{n})$. This implies that the state diagram with $2^{n}$ states (vertices) has exactly $2^{n}$ edges. A nonsingular feedback shift register has a feedback function whose state diagram contains only cycles. Similarly, a factor in a graph is a set of vertex-disjoint cycles that contain all the vertices in the graph. A factor in $G_{n}$ is associated with a state diagram of a nonsingular feedback shift register. These are the only FSRns which are of interest and hence in the rest of the paper all FSRns will always be nonsingular. We can easily verify if a feedback shift register is nonsingular by the following lemma.
Lemma 2
An $\textup{FSR}_{n}$ is nonsingular if and only if for each $x_{i}\in\{0,1\}$, $2\leq i\leq n$, we have
$f(0,x_{2},\ldots,x_{n})\neq f(1,x_{2},\ldots,x_{n})$
Corollary 1
An $\textup{FSR}_{n}$ is nonsingular if and only if for each $x_{i}\in\{0,1\}$, $2\leq i\leq n$, we have
$f(x_{1},x_{2},\ldots,x_{n})=x_{1}+g(x_{2},\ldots,x_{n})~{},$
where $g(x_{2},\ldots,x_{n})$ is any Boolean function with $n-1$ variables.
A nonsingular linear feedback shift register of order $n$ (an LFSRn in short) over $\mathbb{F}_{q}$ is an FSRn whose feedback function $f$ is linear, i.e.,
$x_{n+1}=f(x_{1},x_{2},\ldots,x_{n})=x_{1}+\sum_{i=2}^{n}c_{i}x_{i},~{}~{}c_{i}\in\mathbb{F}_{2},~{}~{}1\leq i\leq n~{}.$
The class of linear shift registers is the most important one, but most of the shift registers are nonlinear. While, by Corollary 1 there are $2^{2^{n-1}}$ nonsingular feedback shift registers of order $n$, the definition of LFSRn implies that there are only $2^{n-1}$ nonsingular LFSRns. Some functions of nonlinear shift register are very similar to the functions of the linear ones and these will be of special interest. Four feedback functions will be defined now.
The first one is called the PCRn for the pure cycling register of order $n$, and its linear feedback function is
$f(x_{1},x_{2},\ldots,x_{n})=x_{1}~{}.$
The second FSRn is called the CCRn for the complemented cycling register of order $n$, and its nonlinear feedback function is
@@
The third FSRn, which will be considered, is called the PSRn for the pure summing register of order $n$. Its linear feedback function is
$f(x_{1},x_{2},\ldots,x_{n})=\sum_{i=1}^{n}x_{i}~{}.$
The fourth FSRn is called the CSRn for the complemented summing register of order $n$, and its nonlinear feedback function is
@@
The length of the cycles in the state diagrams of these shift-registers are of special interest. The length of the cycle in the state diagram is the period of the associated sequence. The length of the sequence is the number of symbols written in the sequence. Hence, the length of the sequence $S=[s_{0}s_{1}\cdots s_{n-1}]$ is $n$. The period $\pi(S)$ of the cyclic sequence ${\cal S}=[s_{0}s_{1}\cdots s_{n-1}]$, is the least positive integer $\pi$ such that $s_{i}=s_{\pi+i}$, for each $10\leq i\leq n-$ 1, where indices are taken modulo $n$, when the sequence is cyclic. For a cyclic sequence ${\cal S}=[s_{0}s_{1}\cdots s_{n-1}]$ any integer $\pi$, $1\leq\pi\leq n$, such that $\pi$ divides $n$ and $s_{i}=s_{\pi+i}$ for $10\leq i\leq n-$ 1 is a period of $S$, but the period is the smallest such integer. Two cyclic sequences are said to be equivalent if one is a cyclic shift of the other. Such sequences are usually considered to be the same unless the starting point of the cyclic sequence is determined. Similar definitions for the length and the period are applied on arrays, The length and the period of sequences generated from shift registers were considered in [10, 20]. These comprehensive books, on shift-register sequences, the de Bruijn graph and its sequences, also consider the length of cycles in any state diagram of linear shift register. A linear shift register can be represented and analyzed by polynomials and it will be discussed also in Section II-B. A cycle of the state diagram can be represented by a sequence whose length equal to the length of the cycle. The sequence is generated from the first bits of the vertices in the cycle keeping their order as the order of the vertices in the cycle.
Example 1
The cyclic sequence
$[010010010]$
is of length 9, but its period is 3. It also has periods 6 and 9.
The cyclic sequence
$[00101]$
is of length 5 and its period is 5.
The two sequences
$[0011101111001]~{}~{}~{}\text{and}~{}~{}~{}[1011110010011]$
are equivalent sequences as one is a cyclic shift of the other, i.e.,
$[0011101111001]\simeq[1011110010011].$
$\blacksquare$
@@
The companion of a vertex (state) $x=(x_{1},x_{2},\ldots,x_{n-1},x_{n})$ is the vertex $x^{\prime}=(x_{1},x_{2},\ldots,x_{n-1},\bar{x}_{n})$, where $\bar{b}$ is the binary complement of $x$. The complement of a state $x=(x_{1},x_{2},\ldots,x_{n-1},x_{n})$ is the state $\bar{x}=(\bar{x}_{1},\bar{x}_{2},\ldots,\bar{x}_{n-1},\bar{x}_{n})$. A cycle (sequence) ${\cal S}=[s_{1},s_{2},\ldots,s_{k}]$ is called self-dual if the complement of ${\cal S}$, ${\bar{{\cal S}}=[\bar{s}_{1},\bar{s}_{2},\ldots,\bar{s}_{k}]}$ is equivalent to ${\cal S}$. A sequence ${\cal S}$ is self-dual if and only if it can has the form ${\cal S}=[X~{}\bar{X}]$.
@@
Two cycles ${\cal C}_{1}$ and ${\cal C}_{2}$ with a state $x$ on ${\cal C}_{1}$ and its companion $x^{\prime}$ on ${\cal C}_{2}$ are merged into one cycle ${\cal C}$ when the predecessors of $x$ and $x^{\prime}$, respectively, $z$ on ${\cal C}_{1}$ and $y$ on ${\cal C}_{2}$, respectively, are interchanged. This operation is called the merge-or-split method since it either merges two cycles into one cycle or it splits one cycle into two cycles. The two possible scenarios of the merge-or-split method are depicted in Fig. 2.

### v2

@@ Abstract
A de Bruijn array code is a set of $r\times s$ binary doubly-periodic arrays such that each binary $n\times m$ matrix is contained exactly once as a window in one of the arrays. Such a set of arrays can be viewed as a two-dimensional generalization of a perfect factor in the de Bruijn graph. Necessary conditions for the existence of such codes are given. Several direct constructions and recursive constructions for such arrays are given. A framework for a theory of two-dimensional feedback shift registers which is akin to (one-dimensional) feedback shift registers is suggested in the process.
@@ I Introduction
Generalizations of one-dimensional sequences and codes to higher dimensions are quite natural from both theoretical and practical points of view. Such generalizations were considered for various structures such as error-correcting codes [1, 48], burst-correcting codes [2, 3, 15, 16], constrained codes [51, 55], and de Bruijn sequences [8, 35, 43]. This paper considers a generalization of one-dimensional sequences with a window property to two-dimensional arrays with a window property. For simplicity in this paper only binary arrays and usually binary sequences are considered, although some of the results can be generalized to any alphabet size and some sequences over larger alphabets can be used to construct binary arrays.
@@
A de Bruijn array (often known as a perfect map) is an $r\times s$ doubly-periodic array (cyclic horizontally and vertically like in a torus), such that each $n\times m$ binary matrix appears exactly once as a window in the array.
Perfect maps were first presented by Reed and Stewart [47] and later in Gordon [21] and Clapham [6]. A very simple construction based on a span $n$ de Bruijn sequence and a de Bruijn sequence over an alphabet of size $2^{n}$ was presented by Ma [32]. This construction will be modified later in one of our constructions. Fan, Fan, Ma, and Siu [17] gave a two-dimensional graphical representation for perfect maps. They also presented two recursive constructions for such de Bruijn arrays. An important construction for perfect maps and another graphical representation was given in [8]. Finally, Paterson [43] proved that the necessary conditions for the existence of de Bruijn arrays are also sufficient.
These arrays have found a variety of applications. They were used in pattern recognition for structured light systems as described in Geng [19], Lin, Nie, and Song [30], Morano, Ozturk, Conn, Dubin, Zietz, and Nissanov [41], Salvi, Fernandez, Pribanic, and Llado [49], and Salvi, Pagès, and Batlle [50]. They are also used in transferring planar surface into a sensitive touch screen display, see Dai and Chung [7], in camera localization as described by Szentandrasi, Zachariás̆, Havel, Herout, Dubska, and Kajan [54], and in one-shot shape acquisition, e.g., Pagès, Salvi, Collewet, and Forest [42]. Finally, they can be applied to surface measurements as described in Kiyasu, Hoshino, Yano, and Fujimura [27] and in Spoelder, Vos, Petriu, and Groen [53], and also in coded aperture imaging as was presented for example in Gottesman and Fenimore [23].
The de Bruijn arrays have been extensively covered in the literature. Further work on this topic and related structures was done by [24, 25, 26, 31, 33, 34, 35, 36, 37, 38]
The goal of the current paper is to extend the results on de Bruijn arrays into de Bruijn array codes defined as follows.
@@
A de Bruijn array code is a set of $r\times s$ doubly-periodic arrays, such that each nonzero $n\times m$ matrix appears exactly once as a window in one of the arrays. Such a set of arrays will be referred to as $(r,s;n,m)$-DBAC. The size $\Delta$ of an array code is the number of arrays (codewords) in the code.
@@
Lemma 1 implies that $r$, $s$, and $\Delta$, are powers of two. Henceforth, we use $2^{k}$ instead of $r$ and $2^{t}$ instead of $s$. One of the goals in the research on DBACs is to prove that the necessary conditions of Lemma 1 are sufficient. In the discussion given here it will also be shown how the definition of de Bruijn array codes generalizes some one-dimensional definitions. For a given $(2^{k},2^{t};n,m)$-BBAC we can always assume that $2\leq k,n$ and $1\leq m,t$. Further restriction associated with some constructions will be given later.
@@
The rest of this paper is organized as follows. In Section II we present the basic definitions and theory for the arrays considered here. In particular the section introduces some theory of shift-register sequences and factors (state diagrams) of the de Bruijn graph. Many constructions are presented in the paper, some of these constructions are direct constructions and some of them are recursive constructions. Section III is devoted to a few direct constructions. Recursive constructions will be presented in Section IV. Perfect factors have to be used carefully for these constructions and the chosen arrays on which the constructions are applied should be also chosen in an appropriate way. In Section V another technique to generate $(2^{k},2^{t};n,m)$-DBACs by joining codewords (cycles) of a $(2^{k},2^{t^{\prime}};n,m)$-DBAC with shorter codewords, i.e., $t^{\prime}<t$, is presented. This technique can be applied if all the codewords are cycles in the same state diagram of the two-dimensional generalization de Bruijn graph. Section VI contains an analysis of the constructions and their input. It also summarizes of some parameters of $(2^{k},2^{t};n,m)$-DBACs obtained here. Section VII includes conclusions and future work.
@@ II-A Shift-register sequences
A feedback shift register of order $n$ (an FSRn in short) has $2^{n}$ states, represented by the set of $2^{n}$ binary words of length $n$. A register has $n$ cells, which are binary storage elements, where each cell stores at each stage one of the bits of the current state $x=(x_{1},x_{2},\ldots,x_{n})$. An FSRn is depicted in Fig. 1.
@@
Shift registers and their sequences have been extensively studied and three comprehensive books on them are [10, 20, 22]. The material in this section is presented in these books. If the word $(x_{1},x_{2},\ldots,x_{n})$ is the state of the FSRn, then $x_{i}$ is stored in the $i$-th cell of the FSRn. The $n$ cells are connected to another logic element which computes a Boolean feedback function $f(x_{1},x_{2},\ldots,x_{n})$. At periodic intervals, controlled by a master clock, $x_{2}$ is transferred to $x_{1}$, $x_{3}$ to $x_{2}$, and so on until $x_{n}$ is transferred to $x_{n-1}$. The value of the feedback function is transferred to $x_{n}$ and hence it is common to write $x_{n+1}=f(x_{1},x_{2},\ldots,x_{n})$. The register starts with an initial state $(a_{1},a_{2},\ldots,a_{n})$, where $a_{i}$, $1\leq i\leq n$, is the initial value stored in the $i$-th cell. There are $2^{n}$ possible distinct values for $x_{1},x_{2},\ldots,x_{n}$, i.e., there are $2^{n}$ distinct states and each one can have a value of either 0 or 1. Hence, there are $2^{2^{n}}$ different FSRn, but not all of these functions are of interest here.
Each FSRn has a state diagram, a graph with $2^{n}$ vertices (the states of the FSRn). Given an FSRn with feedback function $f(x_{1},x_{2},\ldots,x_{n})$, the vertex $(x_{1},x_{2},\ldots,x_{n})$ in the associated state diagram has an edge to the vertex $(x_{2},\ldots,x_{n},x_{n+1})$ if $x_{n+1}=f(x_{1},x_{2},\ldots,x_{n})$. This implies that the state diagram with $2^{n}$ states (vertices) has exactly $2^{n}$ edges. A nonsingular feedback shift register has a feedback function whose state diagram contains only cycles. A factor in a graph is a set of vertex-disjoint cycles that contain all the vertices in the graph; a factor in $G_{n}$ is associated with a state diagram of a nonsingular feedback shift register. These are the only FSRns which are of interest here. Two classes of nonlinear FSRs are of interest here.
The first class denoted by CCRn for the complemented cycling register of order $n$, has the feedback function
@@
The second class denoted by CSRn for the complemented summing register of order $n$, has the feedback function
@@
The length of the cycles in the state diagrams of these shift-registers are of special interest. The length of a cycle in the state diagram is the period of the associated sequence. A cyclic sequence will be denoted by $[s_{0}s_{1}\cdots s_{n-1}]$ and an acyclic sequence (a word) by $(s_{0}s_{1}\cdots s_{n-1})$. A doubly-periodic array is cyclic both horizontally and vertically. The length of the sequence is the number of symbols in the sequence. Hence, the length of the sequence $S=[s_{0}s_{1}\cdots s_{n-1}]$ is $n$. The period $\pi(S)$ of the cyclic sequence ${\cal S}=[s_{0}s_{1}\cdots s_{n-1}]$, is the least positive integer $\pi$ such that $s_{i}=s_{\pi+i}$, for each $10\leq i\leq n-$ 1, where indices are taken modulo $n$, when the sequence is cyclic. Two cyclic sequences are said to be equivalent if one is a cyclic shift of the other. Such sequences are usually considered to be the same unless the starting point of the cyclic sequence is determined. Similar definitions for the length and the period are applied to arrays. A cycle of the state diagram can be represented by a sequence with length equal to the length of the cycle. The sequence is generated from the first bits of the vertices in the cycle keeping their order as the order of the vertices in the cycle. The period of a $2^{k}\times 2^{t}$ doubly-periodic array can be less than $2^{k}$ vertically and less than $2^{t}$ horizontally since doubly-periodic indicates only that the array is cyclic in both directions. However, this scenario will not happened in the arrays which will be discussed in the paper.
@@
The companion of a vertex (state) $x=(x_{1},x_{2},\ldots,x_{n-1},x_{n})$ is the vertex $x^{\prime}=(x_{1},x_{2},\ldots,x_{n-1},\bar{x}_{n})$, where $\bar{b}$ is the binary complement of $x$. The complement of a state $x=(x_{1},x_{2},\ldots,x_{n-1},x_{n})$ is the state $\bar{x}=(\bar{x}_{1},\bar{x}_{2},\ldots,\bar{x}_{n-1},\bar{x}_{n})$. A cycle (sequence) ${\cal S}=[s_{1},s_{2},\ldots,s_{k}]$ is called self-dual if the complement of ${\cal S}$, ${\bar{{\cal S}}=[\bar{s}_{1},\bar{s}_{2},\ldots,\bar{s}_{k}]}$ is equivalent to ${\cal S}$. A sequence ${\cal S}$ is self-dual if and only if it has the form ${\cal S}=[X~{}\bar{X}]$ , where the length of ${\cal S}$ is also its period.
@@
Two cycles ${\cal C}_{1}$ and ${\cal C}_{2}$ with a state $x$ in ${\cal C}_{1}$ and its companion $x^{\prime}$ in ${\cal C}_{2}$ are merged into one cycle ${\cal C}$ when the predecessors of $x$ and $x^{\prime}$, $z$ on ${\cal C}_{1}$ and $y$ on ${\cal C}_{2}$, respectively, are interchanged. This operation is called the merge-or-split method since it either merges two cycles into one cycle or splits one cycle into two cycles. The two possible scenarios of the merge-or-split method are depicted in Fig. 2.